# SoundnessBench: A Soundness Benchmark for Neural Network Verifiers

### v2

@@ 4.1 Implementation Details
For every model in the benchmark, we apply our aforementioned training framework with the following hyperparameters: (1) Number of training epochs is set to 5000; (2) Adam optimizer (Kingma and Ba, 2015) with a cyclic learning rate schedule is used, where the learning rate gradually increases from 0 to a peak value during the first half of the training, and then it gradually decreases to 0 during the second half of the training, with the peak learning rate set to $0.001$; (3) The threshold in the margin objective defined in Eq. (9) is set to $\lambda=0.01$; (4) The size of the perturbation sliding window used in Eq. (10) is set to $w=300$; (5) We use the PGD attack (Madry et al., 2018) for the adversarial training with up to 150 restarts (number of different random initialization for the PGD attack) and 150 attack steps (number of PGD steps in the attack).
@@ 4.2 Training NNs and Building the Benchmark
To make sure that our pre-defined counterexamples are truly hidden against adversarial attacks that try to find counterexamples, we evaluate them using strong adversarial attacks. Specifically, we apply PGD attack with strong attack parameters, $1000$ restarts and $5000$ PGD steps, and also AutoAttack (Croce and Hein, 2020) which combines multiple adversarial attack strategies. For the pre-defined counterexamples, only those passing all of these attacks are considered to be hidden counterexamples.
@@ 4.3 Testing NN Verifiers
We test the soundness of four representative NN verifiers, including $\alpha,\!\beta$-CROWN [^1] https://github.com/Verified-Intelligence/alpha-beta-CROWN_vnncomp2024 (Zhang et al., 2018; Xu et al., 2020, 2021; Wang et al., 2021; Zhang et al., 2022; Shi et al., 2025), NeuralSAT [^2] https://github.com/dynaroars/neuralsat/commit/8c6f006c1acb293ce3f5a995935fd7bc1a7851f5 (Duong et al., 2024), PyRAT [^3] https://git.frama-c.com/pub/pyrat/-/tree/vnncomp2024 (Lemesle et al., 2024), and Marabou (Katz et al., 2019; Wu et al., 2024). For $\alpha,\!\beta$-CROWN and NeuralSAT, we consider two distinct splitting strategies in their branch-and-bound for the verification, activation splitting and input splitting, which conduct branch-and-bound on the activation space and input space, respectively. For Marabou, we include two versions: Marabou 2023444https://github.com/wu-haoze/Marabou/tree/vnn-comp-23 and Marabou 2024555https://github.com/wu-haoze/Marabou/tree/vnn-comp-24, which were used in VNN-COMP 2023 and VNN-COMP 2024, respectively. We include Marabou 2023 because it supported ViT models, while Marabou 2024 does not. To streamline our experiment tables, we define abbreviations for different versions of verifiers, as listed in Table 3. A timeout of 100 seconds is applied on all the verification experiments, similar to what many existing benchmarks use (Bak et al., 2021; Müller et al., 2022; Brix et al., 2023a, 2024). If a verifier fails to output a result for an instance within the time limit, the instance is claimed as neither verified nor falsified.
@@
Table 4: Proportion of “verified” unverifiable instances. Any verified instance indicates a bug in the NN verifier. “-” indicates the model is not supported by the verifier. The number of unverifiable instances used in each setting corresponds to “hidden CEX” in LABEL:tab:training_results (only unverifiable instances with hidden counterexamples not found by adversarial attacks are kept, as mentioned in Section˜4.2).
Model Verified (%) Name $\epsilon$ Input size ABC-A ABC-I NSAT-A NSAT-I PyRAT MB-23 MB-24 CNN AvgPool 0.2 $1\times 5\times 5$ 0 0 0 0 0 0 0 CNN AvgPool 0.2 $3\times 5\times 5$ 100 100 100 100 0 0 0 CNN AvgPool 0.5 $1\times 5\times 5$ 100 66.7 100 33.3 0 0 0 CNN AvgPool 0.5 $3\times 5\times 5$ 0 0 0 0 0 0 0 CNN Tanh 0.2 $1\times 5\times 5$ 0 0 0 0 0 0 0 CNN Tanh 0.2 $3\times 5\times 5$ 0 0 0 0 0 100 0 CNN Sigmoid 0.2 $1\times 5\times 5$ 0 0 0 0 0 0 0 CNN Sigmoid 0.2 $3\times 5\times 5$ 0 0 0 0 0 0 0 ViT 0.2 $1\times 5\times 5$ 0 0 0 0 - 90 - ViT 0.2 $3\times 5\times 5$ 0 0 0 0 - 100 -
@@
    Our benchmark successfully identified bugs in three well-established verifiers: including $\alpha,\!\beta$-CROWN, NeuralSAT, and Marabou 2023. LABEL:tab:verified_unv shows the proportion of “verified” instances claimed by the verifiers on unverifiable instances. For the entries with nonzero values, the results show that these verifiers’ claims are unsound and the verifiers contain bugs in these cases.
@@
Model Architecture ABC-A ABC-I NSAT-A NSAT-I PyRAT MB-23 MB-24 CNN 1 Conv 45.0 35.0 47.5 35.0 25.0 7.5 45.0 CNN 2 Conv 32.5 27.5 32.5 27.5 12.5 5.0 22.5 CNN 3 Conv 17.5 20.0 20.0 20.0 12.5 0 20.0 CNN AvgPool 75.0 70.0 75.0 70.0 0 0 0 CNN Tanh 0 0 0 0 0 50.0 0 CNN Sigmoid 10.0 10.0 10.0 10.0 10.0 10.0 5.0 ViT 0 0 0 0 - 100.0 - MLP 4 Hidden 55.0 75.0 55.0 85.0 45.0 0 45.0 MLP 5 Hidden 45.0 65.0 50.0 75.0 15.0 20.0 15.0
@@
For the bugs on CNN AvgPool models, triggered for $\alpha,\!\beta$-CROWN and NeuralSAT, we find that they are rooted in the auto_LiRPA package (Xu et al., 2020) leveraged by both $\alpha,\!\beta$-CROWN and NeuralSAT in their bound computation. Specifically, we find that auto_LiRPA previously assumed that “strides” and “kernel size” are equal in AvgPool layers and silently produced unsound results when their assumption does not hold in our models. Counterexamples provided by SoundnessBench successfully led to the discovery of the bug in auto_LiRPA. We shared our test cases that helped the developers of auto_LiRPA fix the bug. The fix has been reflected in the auto_LiRPA package and adopted by both $\alpha,\!\beta$-CROWN and NeuralSAT in their latest version. For the bugs of Marabou on Tanh and ViT, the authors of Marabou conjectured that there might be bugs in the interaction with the quadratic programming engine in Gurobi, but the specific cause remains unknown.
@@ 4.4 Ablation study
Table 7: Average number of hidden counterexamples (up to 10) with or without the margin objective, evaluated across 5 different random seeds.
Training Settings Margin objective setting Model Name $\epsilon$ Input size Disabled Enabled CNN 1 Conv 0.2 $1\times 5\times 5$ 1.2 6.2 CNN 1 Conv 0.2 $3\times 5\times 5$ 1.2 10.0 CNN 1 Conv 0.5 $1\times 5\times 5$ 0.4 7.4 CNN 1 Conv 0.5 $3\times 5\times 5$ 0 5.8
Table 8: Average number of hidden counterexamples (up to 10) with various perturbation sliding window sizes, evaluated across 5 different random seeds.
Training Settings Perturbation window size Model Name $\epsilon$ Input size 1 10 100 300 CNN 1 Conv 0.2 $1\times 5\times 5$ 0 0 4.0 6.2 CNN 1 Conv 0.2 $3\times 5\times 5$ 0 0 8.0 10.0 CNN 1 Conv 0.5 $1\times 5\times 5$ 0 0 3.6 7.4 CNN 1 Conv 0.5 $3\times 5\times 5$ 0 0 6.6 5.8
@@
LABEL:tab:margin_objective compares the results with or without margin objective. Models trained with the margin objective consistently outperform their counterparts trained without the margin objective, demonstrating the effectiveness of the margin objective in producing many more hidden counterexamples. LABEL:tab:perturbation_window compares the results when using four different maximum sliding window sizes. When the window size is set to 1, the training is equivalent to the original adversarial training, and no hidden counterexample can be produced. The best results are achieved by setting the window size to $100$ in one case and to $300$ in the other three cases. This demonstrates the effectiveness of our perturbation sliding window.
5 Conclusion
In this paper, we present SoundnessBench, a benchmark with hidden counterexamples for testing the soundness of NN verifiers. Using a two-objective training framework with two training techniques, margin objective and perturbation sliding window, we incorporate a total of 26 models across nine distinct NN architectures, 206 unverifiable instances with hidden counterexamples, and 260 regular instances. Notably, SoundnessBench revealed bugs in three well-established NN verifiers: $\alpha,\!\beta$-CROWN, NeuralSAT, and Marabou 2023. Our benchmark has helped the authors of auto_LiRPA, which is used in both $\alpha,\!\beta$-CROWNand NeuralSAT, fix the bug we identified. We believe SoundnessBench and our proposed methodology can serve as a valuable resource for developers building more reliable NN verifiers and more trustworthy artificial intelligence, thereby broadly benefiting safety-critical domains utilizing NN verifiers in the future.
@@ 5 Conclusion
This work has several limitations. First, our models and data are relatively small-scale, as discussed in Section˜3.5, but they can still be useful for testing the soundness of NN verifiers as demonstrated. Second, we mainly use PGD and AutoAttack, which are among the strongest and most commonly used attacks in verifiers (such as $\alpha,\!\beta$-CROWN) and the literature. While PGD and AutoAttack cannot cover new attacks that may appear in the future, it is only feasible for us to adopt currently existing attacks, and future work may consider new attacks as we discuss below.
@@
Our benchmark can be extended to consider new settings and new attacks that may emerge. We suggest closely monitoring the development of various verifiers and performance benchmarks to maintain our benchmark: (1) Results should be updated as new releases of verifiers appear; (2) If a new attack is introduced in a verifier, we need to test the resilience of our currently hidden counterexamples against such attacks. If our counterexamples are no longer hidden, we need to adapt our training to incorporate such stronger attacks by replacing the PGD attack with the new attack algorithm, as the attack component in our training pipeline is independent of the other components; (3) If a new model architecture or data setting becomes popular in the NN verification community and is supported by multiple verifiers, we can extend our benchmark to support such settings. We welcome contributions from the community to help maintain and expand SoundnessBench.
Appendix A Synthetic Bugs

### v3

@@ 4.1 Implementation Details
For every model in the benchmark, we apply our aforementioned training framework with the following hyperparameters: (1) Number of training epochs is set to 5000; (2) Adam optimizer (kingma2014adam) with a cyclic learning rate schedule is used, where the learning rate gradually increases from 0 to a peak value during the first half of the training, and then it gradually decreases to 0 during the second half of the training, with the peak learning rate set to $0.001$; (3) The threshold in the margin objective defined in Eq. (10) is set to $\lambda=0.01$; (4) The size of the perturbation sliding window used in Eq. (11) is set to $w=300$; (5) We use the PGD attack (madry2017towards) for the adversarial training with up to 150 restarts (number of different random initialization for the PGD attack) and 150 attack steps.
@@ 4.2 Training NNs and Building the Benchmark
To make sure that our pre-defined counterexamples are truly hidden against adversarial attacks that try to find counterexamples, we evaluate them using strong adversarial attacks. Specifically, we apply PGD attack with strong attack parameters, $1000$ restarts and $5000$ PGD steps, and also AutoAttack (croce2020reliable) which combines multiple adversarial attack strategies. For the pre-defined counterexamples, only those passing all of these attacks are considered to be hidden counterexamples.
@@ 4.3 Testing NN Verifiers
We test the soundness of four representative NN verifiers, including $\alpha,\!\beta$-CROWN [^1] https://github.com/Verified-Intelligence/alpha-beta-CROWN_vnncomp2024 (zhang2018efficient; xu2020automatic; xu2020fast; wang2021beta; zhang2022general; shi2024genbab), NeuralSAT [^2] https://github.com/dynaroars/neuralsat/commit/8c6f006c1acb293ce3f5a995935fd7bc1a7851f5 (duong2024harnessing), PyRAT [^3] https://git.frama-c.com/pub/pyrat/-/tree/vnncomp2024 (lemesle2024neural), and Marabou (katz2019marabou; wu2024marabou). For $\alpha,\!\beta$-CROWN and NeuralSAT, we consider two distinct splitting strategies in their branch-and-bound for the verification, activation splitting and input splitting, which conduct branch-and-bound on the activation space and input space, respectively. For Marabou, we include two versions: Marabou 2023444https://github.com/wu-haoze/Marabou/tree/vnn-comp-23 and Marabou 2024555https://github.com/wu-haoze/Marabou/tree/vnn-comp-24, which were used in VNN-COMP 2023 and VNN-COMP 2024, respectively. We include Marabou 2023 because it supported ViT models, while Marabou 2024 does not. To streamline our experiment tables, we define abbreviations for different versions of verifiers, as listed in Table 3. A timeout of 100 seconds is applied on all the verification experiments, similar to what many existing benchmarks use (bak2021second; vnn2022; vnn2023; brix2024fifth). If a verifier fails to output a result for an instance within the time limit, the instance is claimed as neither verified nor falsified.
@@
Table 4: Proportion of “verified” unverifiable instances. It is a measurement of unsoundness of verifiers. Any verified instance indicates a bug in the NN verifier. “-” indicates the model is not supported by the verifier. The number of unverifiable instances used in each setting corresponds to “hidden CEX” in LABEL:tab:training_results (only unverifiable instances with hidden counterexamples not found by adversarial attacks are kept, as mentioned in Section˜4.2).
Model Verified (%) Name $\epsilon$ Input size ABC-A ABC-I NSAT-A NSAT-I PyRAT MB-23 MB-24 CNN AvgPool 0.2 $1\times 5\times 5$ 0 0 0 0 0 - - CNN AvgPool 0.2 $3\times 5\times 5$ 100 100 100 100 0 - - CNN AvgPool 0.5 $1\times 5\times 5$ 100 66.7 100 33.3 0 - - CNN AvgPool 0.5 $3\times 5\times 5$ 0 0 0 0 0 - - CNN Tanh 0.2 $1\times 5\times 5$ 0 0 0 0 0 0 0 CNN Tanh 0.2 $3\times 5\times 5$ 0 0 0 0 0 100 0 CNN Sigmoid 0.2 $1\times 5\times 5$ 0 0 0 0 0 0 0 CNN Sigmoid 0.2 $3\times 5\times 5$ 0 0 0 0 0 0 0 ViT 0.2 $1\times 5\times 5$ 0 0 0 0 - 90 - ViT 0.2 $3\times 5\times 5$ 0 0 0 0 - 100 -
@@
    Our benchmark successfully identified bugs in three well-established verifiers: including $\alpha,\!\beta$-CROWN, NeuralSAT, and Marabou 2023. LABEL:tab:verified_unv shows the proportion of “verified” instances claimed by the verifiers on unverifiable instances, measuring the unsoundness of verifiers. For the entries with nonzero values, the results show that these verifiers’ claims are unsound and the verifiers contain bugs in these cases.
@@
Model Architecture ABC-A ABC-I NSAT-A NSAT-I PyRAT MB-23 MB-24 CNN 1 Conv 45.0 35.0 47.5 35.0 25.0 7.5 45.0 CNN 2 Conv 32.5 27.5 32.5 27.5 12.5 5.0 22.5 CNN 3 Conv 17.5 20.0 20.0 20.0 12.5 0 20.0 CNN AvgPool 75.0 70.0 75.0 70.0 0 - - CNN Tanh 0 0 0 0 0 50.0 0 CNN Sigmoid 10.0 10.0 10.0 10.0 10.0 10.0 5.0 ViT 0 0 0 0 - 100.0 - MLP 4 Hidden 55.0 75.0 55.0 85.0 45.0 0 45.0 MLP 5 Hidden 45.0 65.0 50.0 75.0 15.0 20.0 15.0
@@
For the bugs on CNN AvgPool models, triggered for $\alpha,\!\beta$-CROWN and NeuralSAT, we find that they are rooted in the auto_LiRPA package (xu2020automatic) leveraged by both $\alpha,\!\beta$-CROWN and NeuralSAT in their bound computation. Specifically, we find that auto_LiRPA previously assumed that “strides” and “kernel size” are equal in AvgPool layers and silently produced unsound results when their assumption does not hold in our models. Counterexamples provided by SoundnessBench successfully led to the discovery of the bug in auto_LiRPA. We shared our test cases that helped the developers of auto_LiRPA fix the bug. The fix has been reflected in the auto_LiRPA package and adopted by both $\alpha,\!\beta$-CROWN and NeuralSAT in their latest version. For the bugs of Marabou on Tanh and ViT, the authors of Marabou conjectured that there might be bugs in the interaction with the quadratic programming engine in Gurobi, but the specific cause remains unknown.
@@ 4.4 Ablation study
Table 7: Average number of hidden counterexamples (up to 10) for the ablation study, evaluated across 5 different random seeds.
Training Settings Margin objective Perturbation window size Model $\epsilon$ Input size Disabled Enabled 1 10 100 300 CNN 1 Conv 0.2 $1\times 5\times 5$ 1.2 6.2 0 0 4.0 6.2 CNN 1 Conv 0.2 $3\times 5\times 5$ 1.2 10.0 0 0 8.0 10.0 CNN 1 Conv 0.5 $1\times 5\times 5$ 0.4 7.4 0 0 3.6 7.4 CNN 1 Conv 0.5 $3\times 5\times 5$ 0 5.8 0 0 6.6 5.8
@@
Table˜7 shows the results. Models trained with the margin objective consistently outperform their counterparts trained without the margin objective, demonstrating the effectiveness of the margin objective in producing many more hidden counterexamples. When the window size is set to 1, the training is equivalent to the original adversarial training, and no hidden counterexample can be produced. The best results are achieved by setting the window size to $100$ in one case and to $300$ in the other three cases. This demonstrates the effectiveness of our perturbation sliding window.
5 Conclusion and Discussion
We present SoundnessBench, a benchmark with hidden counterexamples for testing the soundness of NN verifiers. Through a two-objective training framework with two training techniques, margin objective and perturbation sliding window, we have created a benchmark incorporating 26 models across nine distinct NN architectures, 206 unverifiable instances with hidden counterexamples, and 260 regular instances. Notably, SoundnessBench revealed bugs in three well-established NN verifiers: $\alpha,\!\beta$-CROWN, NeuralSAT, and Marabou 2023. Our benchmark has helped the authors of auto_LiRPA, which is used in both $\alpha,\!\beta$-CROWN and NeuralSAT, fix the bug we identified. We believe SoundnessBench and our proposed methodology can serve as a valuable resource for developers building more reliable NN verifiers and more trustworthy artificial intelligence, thus broadly benefiting safety-critical domains that utilize NN verifiers in the future.
@@ 5 Conclusion and Discussion
This work has several limitations. First, our models and data are relatively small-scale, as discussed in Section˜3.5, but they remain useful for testing the soundness of NN verifiers as our results demonstrate. Second, we mainly use PGD and AutoAttack, which are among the strongest and most commonly used attacks both in verifiers (such as $\alpha,\!\beta$-CROWN) and in the literature. While PGD and AutoAttack cannot cover new attacks that may appear in the future, it is only feasible for us to adopt currently existing attacks, and future work may consider new attacks as we discuss below.
@@
Our benchmark can be extended to consider new settings and new attacks that may emerge. We suggest closely monitoring the development of various verifiers and performance benchmarks to maintain our benchmark: (1) Results should be updated as new releases of verifiers appear; (2) If a new attack is introduced in a verifier, we need to test the resilience of our currently hidden counterexamples against such attacks. If our counterexamples are no longer hidden, we should adapt our training to incorporate such stronger attacks by replacing the PGD attack with the new attack algorithm, as the attack component in our training pipeline is independent of the other components; (3) If a new model architecture or data setting becomes popular in the NN verification community and is supported by multiple verifiers, we can extend our benchmark to support such settings. We welcome contributions from the community to help maintain and expand SoundnessBench.
Increasing the soundness and reliability of NN verifiers.
Beyond the development of novel and advanced NN verification algorithms, the correct implementation of algorithms is critical to maintaining the soundness of verifiers, which can be challenging to validate systematically. Most NN verification papers and VNN-COMP primarily focus on the performance of NN verifiers, while the reliability of NN verifiers is currently understudied. We thus encourage the community to dedicate more effort to testing the soundness of NN verifiers. Specifically, we recommend that researchers: (1) stress-test new NN verifiers with soundness benchmarks and tools when new algorithms or implementation modifications are introduced, and develop new soundness benchmarks simultaneously as new settings are studied; (2) incorporate measures and results of soundness testing into research papers and technical reports; and (3) document the assumptions of algorithms and implementations and properly handle unsupported cases. By prioritizing both performance and soundness, the community can build more capable and reliable NN verifiers and verified NN-based models for various safety-critical applications.
Acknowledgments
This project is supported in part by NSF 2048280, 2331966 and ONR N00014-23-1-2300:P00001. Huan Zhang is supported in part by the AI2050 program at Schmidt Sciences (AI2050 Early Career Fellowship) and NSF (IIS-2331967).
Appendix A Illustration for the Training Algorithm
In Algorithm˜1, we illustrate our algorithm for training models with hidden counterexamples, as proposed in Section˜3.4.
Algorithm 1 Training Models with Hidden Counterexamples
Inputs: Dataset $\mathcal{D}=\{({\bm{x}}^{(i)}_{0},y^{(i)})\}_{i=1}^{2n}$; initialized model ${\bm{f}}_{\bm{\theta}}$; attack method $\mathcal{A}$ (i.e., PGD); margin cap $\lambda$; perturbation sliding window size $w$; epochs $T$; optimizer $\mathtt{Opt}$.
Output: Trained model ${\bm{f}}_{\bm{\theta}}$
1:// Initialize perturbation sliding windows
2:for $i=1,\dots,2n$ do
3:  ${\mathbb{W}}^{(i)}\leftarrow\emptyset$
4:end for
5:// Main training loop
6:for $t=1,2,\dots,T$ do
7:  // Generate new adversarial perturbations and update perturbation sliding windows
8:  for $i=1,\dots,2n$ do
9:   ${\bm{\delta}}^{(i)}\leftarrow\mathcal{A}({\bm{f}}_{\bm{\theta}},{\bm{x}}^{(i)}_{0},y^{(i)},\epsilon)$
10:   ${\mathbb{W}}^{(i)}\leftarrow({\mathbb{W}}^{(i)}\cup\{{\bm{\delta}}^{(i)}\})$ and truncate to keep the most recent $w$ elements
11:  end for
12:  // Margin objective for pre-defined counterexamples (Eq. (10))
13:  ${\mathcal{L}}_{\text{CEX}}=\frac{1}{n}\sum_{i=1}^{n}\max\{0,{\bm{f}}_{y^{(\text{unv},i)}}({\bm{x}}^{(\text{unv},i)})-{\bm{f}}_{y_{\text{CEX}}^{(i)}}({\bm{x}}_{\text{CEX}}^{(i)})+\lambda)\}$
14:  // Adversarial training with perturbation sliding windows (Eq. (11))
15:  ${\mathcal{L}}_{\text{adv}}=\frac{1}{2n}\sum_{1\leq i\leq 2n}\frac{1}{|{\mathbb{W}}^{(i)}|}\sum_{{\bm{\delta}}\in{\mathbb{W}}^{(i)}}{\mathcal{L}}_{\text{CE}}({\bm{f}}({\bm{x}}_{0}^{(i)}+{\bm{\delta}}),y^{(i)}).$
16:  // Update the model ${\bm{f}}_{\bm{\theta}}$ with a combined objective (Eq. (12))
17:  ${\mathcal{L}}({\bm{\theta}})={\mathcal{L}}_{\text{CEX}}({\bm{\theta}})+{\mathcal{L}}_{\text{adv}}({\bm{\theta}})$
18:  ${\bm{\theta}}\leftarrow\mathtt{Opt.\,step}({\bm{\theta}},\nabla_{\bm{\theta}}{\mathcal{L}}({\bm{\theta}})$
19:end for
20:return Trained model ${\bm{f}}_{\bm{\theta}}$
Appendix B Comparison between Benchmarks
In Table˜8, we compare our benchmark with existing ones compiled in VNN-COMP 2024 (brix2024fifth). Our benchmark serves a unique purpose for testing the soundness of NN verifiers while others all focus on performance. Meanwhile, our benchmark covers diverse model architectures, while most existing benchmarks focus on one or two model architectures. Additionally, while we focus on relatively small models and data as discussed in Section˜3.5, our scale is still larger than many of the existing benchmarks.
Table 8: Comparison between different benchmarks. Except for our benchmark, others are cited from VNN-COMP 2024 (brix2024fifth).
Benchmark Purpose Model architecture # Params Input size Acas XU Performance MLP 13k 5 CCTSDB Complex 100k 2 cGAN CNN, ViT 500k - 68M 5 cifar100 ResNet 2.5M - 3.8M 3072 Collins Aerospace Complex 1.8M 1.2M Collins RUL CNN CNN 60k - 262k 400 - 800 CORA MLP 575k, 1.1M 784, 3072 Dist Shift MLP 342k - 855k 792 LinearizeNN MLP 203k 4 LSNC Complex 210k 8 Metaroom Conv 466k - 7.4M 5376 ml4acopf Complex 4k - 680k 22 - 402 NN4Sys MLP 33k - 37M 1-308 safeNLP MLP 4k 30 tinyimagenet ResNet 3.6M 9408 TLL Verify Bench Two-Level Lattice NN 17k - 67M 2 Traffic Signs Recognition CNN 905k - 1.7M 2.7k - 12k VGGNet16 CNN 138M 150k ViT ViT 68k - 76k 3072 Yolo ResNet 22k - 37M 1 - 308 SoundnessBench (ours) Soundness MLP, CNN (ReLU, Sigmoid, Tanh, AvgPool), ViT 5k - 3M 25 - 75
Appendix C Synthetic Bugs